OpenStudy (anonymous):

@mathstudent55 Please help.

OpenStudy (mathstudent55):

In order to answer this question, we need to see if the two triangles (the gardens) are congruent. In each triangle, we have a 60-ft side. Then, to the left of each 60-ft side there is an angle. The figure shows the angles to the left are congruent. To the right of the 60-ft side, there is a right angle. All right angles are congruent, so by ASA, the two triangles are congruent.

OpenStudy (mathstudent55):

Now that we know the two triangles are congruent, we can use the info form the two triangles interchangeably.

OpenStudy (mathstudent55):



OpenStudy (anonymous):

Would it be x over 30 x ---- ? 30

OpenStudy (mathstudent55):

The formula for the area of a triangle is \(A = \dfrac{1}{2}bh\) We can use the 60-ft side as the base (b), and the side of length x as the height (h). The area (A) is 300 sq ft.

OpenStudy (mathstudent55):

\(A = \dfrac{1}{2} bh\) \(300 = \dfrac{1}{2} \times 60x\) \(300 = 30x\) \(30 x = 300\) \(x = \dfrac{300}{30} \) \(x = 10\)

OpenStudy (mathstudent55):

Sorry, gtg.

OpenStudy (anonymous):

Thank you ! :D

OpenStudy (mathstudent55):

You're welcome.